OpenStudy (anonymous):

find dy/dx y=(cosx)/(1+sinx) I have it figured out I think, but I just want to get it checked to see if I am approaching the problem the correct way. I will post my answer and how I got it in a minute.

OpenStudy (anonymous):

Did you use quotient rule?

OpenStudy (anonymous):

my answer is \[y'=-1/(1+\sin x)\]

OpenStudy (anonymous):

is that where it is \[(u*dv/dx -v*du/dx)v ^{2}\]? where the numerator is u and denominator is v?

OpenStudy (anonymous):

[(1+sinx)(-sinx) - (cosx)(cosx)]/(1+sinx)^2 [sinx/1+sinx] - [(cosx)^2] / [(1+sinx)^2] is what I got

OpenStudy (anonymous):

\[\frac{u'v+uv'}{v^2}=\frac{sin(x)(1+sin(x))+cos(x)(-cos(x))}{(1+sin(x))^2}\]

OpenStudy (anonymous):

@VeritasVosLiberabit when I factored out, I got (-(sin x +sin^2 x +cos^2 x))/(1+sin x)^2

OpenStudy (anonymous):

we have different signs I could be wrong have to check again

OpenStudy (anonymous):

timmtamm that's it. didn't see you factor out the negative

OpenStudy (anonymous):

the numerator was originally -sin x -sin^2 x -cos ^2 x but I took out the negative to get factor it out to -(sin x +sin^2 x +cos^2 x) to then simplify it out to -(sin x +1) as the numerator

OpenStudy (anonymous):

and then -(sin x +1)/(1+sin x)^2 can then be simplified to -1/(1+sin x) right?